OpenStudy (cutestgirlaround):

Randy measured the length of the bolts that came with his model airplane kit and plotted their lengths, in inches, on the line plot. What is the total length of all the bolts that were inches long? Line plot from 0 to 1, labeled every one-eighth. One x mark is above one-eighth. Two x marks are above one-fourth. Five x marks are above one-half. One x mark is above five-eighths. Three x marks are above three-fourths. Two x marks are above seven-eighths. A. in. B. in. C. in. D. 2 in.

OpenStudy (cutestgirlaround):

A.\[\frac{ 7}{ 16}\] B.\[1\frac{ 4}{ 8 }\] C.\[1\frac{ 6}{ 8}\]

OpenStudy (cutestgirlaround):

@Kitkat77

OpenStudy (cutestgirlaround):

@JoshuaBurt

OpenStudy (cutestgirlaround):

@yamyam70

OpenStudy (cutestgirlaround):

please help its for my sis

OpenStudy (cutestgirlaround):

@imqwerty

OpenStudy (cutestgirlaround):

@Serenity74

OpenStudy (cutestgirlaround):

@pooja195

OpenStudy (yamyam70):

this dang X MARKS!

OpenStudy (yamyam70):

@cutestgirlaround sorry my time is short here.. but here is a resource maybe it will help Goodluck! http://www.cehd.umn.edu/ci/rationalnumberproject/88_3.html

OpenStudy (cutestgirlaround):

does that help

OpenStudy (yamyam70):

yeah complete explanation plus examples..try :)

OpenStudy (yamyam70):

"Identifying fractions on number lines" thats your topic

OpenStudy (cutestgirlaround):

yeah its my sis and she needs help so you know i am trying to help her

OpenStudy (yamyam70):

okay goodluck to you both, a good read :)